ROTATIONAL SPECTRA OF CH_{3}CCH-NH_{3}, NCCCH-NH_{3}, AND NCCCH-OH_{2}

\underline{A.R. Hight Walker}, R. Omron, G.T. Fraser, R.D. Suenram, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899; G. Hilpert, Institut für Angewandte Physik, Universität Bonn, Wegelerstraße 8, D-53115 Bonn, Germany.

Microwave spectra of NCCCH-NH_{3}, CH_{3}CCH-NH_{3}, and NCCCH-OH_{2} have been recorded using a pulsed-nozzle Fourier-transform microwave spectrometer. The complexes, NCCCH-NH_{3} and CH_{3}CCH-NH_{3}, are found to have symmetric-top structures with the acetylenic proton hydrogen-bonded to the nitrogen of the NH_{3}. The data for CH_{3}CCH-NH_{3} are further consistent with free or nearly free internal rotation of the methyl top against the ammonia top. For NCCCH-OH_{2}, the acetylenic proton is hydrogen-bonded to the oxygen of the water. The water complex has a dynamical C_{2v} structure, as evidenced by the presence of two nuclear-spin modifications of the complex. The hydrogen-bond lengths and hydrogen-bond stretching force constants are 2.212 Å and 10.8 N/m, 2.322 Å and 6.0 N/m, and 2.125 Å and 9.6 N/m, for NCCCH-NH_{3}, CH_{3}CCH-NH_{3}, and NCCCH-OH_{2}, respectively. For the cyanoacetylene complexes, these bond lengths and force constants lie between the values for the related hydrogen cyanide and acetylene complexes of NH_{3} and H_{2}O. The NH_{3} bending and weak-bond stretching force constants for CH_{3}CCH-NH_{3} are less than those found in NCCCH-NH_{3}, NCH-NH_{3}, and HCCH-NH_{3}, suggesting that the hydrogen bonding interaction is particularly weak in CH_{3}CCH-NH_{3}. The weakness of the bond is partially a consequence of the orientation of the monomer electric dipole moments in the complex. In CH_{3}CCH-NH_{3} the antialigned monomer dipole moments leads to a repulsive dipole-dipole interaction energy, while in NCH-NH_{3} and NCCCH-NH_{3} the aligned dipoles give an attractive interaction.